# A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections

### v2

@@ Abstract
In the proposed approach, we highlight three major contributions. First, we propose a Vision Transformer (ViT) based method for DOA estimation, which achieves excellent performance in scenarios with low signal-to-noise ratios (SNR) and limited snapshots. Second, we introduce a transfer learning framework that extends deep learning models from ideal simulation scenarios to complex real-world scenarios with array imperfections. By leveraging prior knowledge from ideal simulation data, the proposed transfer learning framework significantly improves deep learning-based DOA estimation performance in the presence of array imperfections, without the need for extensive real-world data. Finally, we incorporate visualization and evaluation metrics to assess the performance of DOA estimation algorithms, which allow for a more thorough evaluation of algorithms and further validate the proposed method. Our code can be accessed at https://github.com/zzb-nice/DOA_est_Master.
Index Terms:
Direction of Arrival (DOA) estimation, Domain Adaptive, Vision Transformer, Data-driven Methods, Deep learning
@@ I INTRODUCTION
Direction of Arrival (DOA) estimation finds extensive applications across various engineering fields, including wireless communications, astronomical observation, radar, and sonar [2, 3, 4, 5, 6], where accurately determining the direction of incoming signals is crucial for efficient operation. Today, as radar technology becomes increasingly accessible for civilian use, DOA estimation algorithms have also been applied in emerging areas such as autonomous driving and drone operating system [7, 8, 9, 10].
With the continuous development of engineering technologies, practical scenarios are placing higher demands on DOA estimation algorithms. For example, in autonomous driving scenarios, there is an increasing need for DOA estimation of fast-moving small targets [7, 11]. In non-cooperative target localization scenarios, it is essential to achieve rapid and precise DOA estimation [12, 13]. These practical scenarios generally require accurate DOA estimation results under conditions of low SNR and small snapshots, along with high real-time performance. Therefore, there is an urgent need for a new algorithm with relatively low computational complexity, one that can achieve strong performance in ideal scenarios while also demonstrate improved performance in low SNR and small snapshots scenarios.
@@
Over the past few decades, extensive research has been conducted on DOA estimation, leading to the development of numerous algorithms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 1, 29, 30, 31, 32, 33, 34, 35]. These methods can be broadly categorized into model-driven and data-driven approaches. Among them, model-driven methods have been more extensively studied due to their well-established theoretical foundations. The core principle of model-driven DOA estimation is to establish a mapping between the directions of incoming signals and the received array data, and then use a series of techniques to accurately invert this mapping based on the received data. Based on different estimation strategies, model-driven approaches can be further classified into several major categories, including beamforming methods [14], subspace methods [15, 16, 17, 18, 19], maximum likelihood methods [20, 21, 22], and compressed sensing (sparsity-inducing) methods [36, 23, 24, 25].
@@
A leading and well-respected family of model-driven approaches is the subspace method, which achieves satisfactory performance in scenarios with high signal-to-noise ratio (SNR) and large snapshots. It maintains relatively low computational complexity when a limited number of array elements is used. However, its performance deteriorates significantly in practical scenarios with low SNR and small snapshots. Moreover, subspace-based methods struggle with color noise, and require the estimated number of signal sources to be smaller than the number of array elements, which limits their applicability. Compressive sensing-based and maximum likelihood-based methods have achieved certain improvements but are limited by real-time performance constraints. In summary, model-driven methods for DOA estimation have reached a performance bottleneck in practical scenarios.
In recent years, the rapid advancement of computational hardware [37] has significantly accelerated the development of deep learning technologies. As data-driven technologies, machine learning and deep learning methods have gained increasing attention in the field of DOA estimation. These methods build complex mapping relationships from input to output based on large-scale datasets. Unlike model-driven methods that rely on predefined signal models, data-driven methods generally do not impose prior assumptions on the input and output. This makes them particularly suitable for practical scenarios where the signal model may be incomplete or inaccurate. Furthermore, during practical implementation, data-driven methods demonstrate better real-time performance due to their fast inference speed. As a result, data-driven methods are considered promising tools for achieving DOA estimation in practical scenarios.
Deep learning methods were first introduced to address the problem of DOA estimation in the 1990s [26, 27]. Following a considerable period of minimal utilization, these approaches have witnessed a substantial resurgence in DOA estimation research during recent years. The investigations conducted in [28, 1] pioneered the implementation of Deep Neural Network (DNN) and Convolutional Neural Network (CNN) for DOA estimation. Subsequently, the method proposed in [29, 30] advanced the field by first partitioning the incoming signal into multiple sub-regions through a network architecture, followed by the deployment of parallel networks across the sub-regions.
@@
Additionally, some methods aim to combine traditional DOA estimation techniques with deep learning methods. Due to their inherent flexibility, deep learning methods can effectively establish diverse mapping relationships. As a result, deep learning techniques are not necessarily limited to direct mapping to the DOA values but can also be applied to map various variables, such as noise-free covariance matrices, signal subspaces, angular intervals of incident signals, and so on. This flexibility makes them well-suited for integration with traditional methods such as subspace techniques and compressed sensing. In 2022, the method proposed in [31, 32] integrated subspace-based methods with deep neural networks. Specifically, a CNN was employed to reconstruct a noise-free covariance matrix, after which a classical subspace-based approach was applied for DOA estimation. Similarly, the approaches in [33, 34] combined deep learning with compressed sensing algorithms. In these works, a DNN is utilized to the extend the covariance matrix to a higher-dimensional space or to compute prior knowledge for the compressed sensing algorithm. Then a sparse reconstruction methods is applied to estimate the incident angles. Furthermore, in 2024, Shmuel et al. [35] introduced SubspaceNet, which directly treated the Root MUSIC algorithm as part of the network. Unlike previous work, SubspaceNet utilized the differentiability of polynomial equation root-finding and matrix decomposition, which allowed for the backpropagation of loss from the Root MUSIC algorithm.
Nevertheless, most of the model-driven and data-driven methods mentioned above do not account for array imperfections in practical systems. The existence of array position biases, amplitude and phase inconsistencies, mutual coupling effects [38], and other array imperfections can affect the DOA estimation accuracy of models. In severe cases, these imperfections can significantly degrade the model’s DOA estimation performance or even cause its DOA estimation to fail.
A series of active calibration [39, 40, 41, 42, 43] and self-calibration [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] methods have been applied to compensate for array imperfections in practical scenarios. Among them, active calibration methods require the setup of known signal sources with predefined DOA directions, which are specifically employed for the compensation of array imperfections. In contrast, self-calibration methods do not require additional auxiliary signal sources and can dynamically calibrate the array imperfections during array system applications, offering significant practical advantages. However, active calibration methods benefit from utilizing known DOA values for array imperfection estimation, leading to higher accuracy. In contrast, self-calibration methods generally estimate array imperfections simultaneously with DOA values, often relying on initialization and struggling to handle more severe array imperfections. As noted in [44], most existing self-calibration methods can perform well only under relatively ideal signal conditions and are ineffective in practical scenarios with low SNR, small snapshots, and spatially proximate sources——precisely the challenges that many DOA estimation algorithms face.
Moreover, existing calibration methods are constrained by numerous limitations. Due to the significant difficulty in directly modeling and analyzing all array imperfections, most existing calibration algorithms can focus on only one of the array imperfections, such as array position biases [39, 42, 43], amplitude-phase inconsistencies [45, 46, 47, 48, 49] or mutual coupling [50, 51, 52, 53, 54]. Only a few studies have explored how to address scenarios involving multiple array imperfections [44, 40, 41]. Some methods are even restricted to calibrating specific array structures such as uniform linearity/circularity array [54, 48, 49, 53], or impose additional assumptions to simplify the model [44]. Therefore, how to effectively address scenarios with multiple array imperfections remains an issue that requires further exploration.
In [29], Liu et al. mentioned that deep learning algorithms exhibit good adaptability to array imperfections, but this adaptability fails in practical scenarios with low SNR and small snapshots. Based on the simulation results, it can be observed that in scenarios with low SNR and small snapshots, deep learning models remain sensitive to array imperfections, requiring the application of calibration models to compensate for array imperfections. However, due to the inherent flexibility of deep learning, it may be possible to directly construct a transfer learning algorithm [55] that transitions from an ideal scenario without array imperfections to a practical scenario with array imperfections, thereby enhancing the model’s adaptability to practical scenarios.
@@
Nonetheless, there is currently a lack of research on the application of transfer learning methods in the field of DOA estimation. To the best of our knowledge, only a few papers [56, 57, 58] have introduced transfer learning methods into DOA estimation. In [56], Wu et al. migrated a network designed for DOA estimation with multiple targets to one for fewer targets, with the primary goal of reducing the computational cost of training. Similarly, in [58], Labbaf et al. migrated a network designed for single-target DOA estimation to multiple targets. In [57], Zhang et al. applied transfer learning to enhance the network’s adaptability to array imperfections, but they only fine-tuned the network from ideal data to array imperfections-affected data, without incorporating more advanced transfer learning techniques.
@@
In this paper, we propose a deep learning-based supervised transfer learning framework for DOA estimation, which aims to mitigate the performance degradation of data-driven approaches in practical scenarios with various array imperfections. In our approach, the data is divided into source and target domains. The source domain data consists of ideal data without array imperfections, while the target domain data represents practical scenarios containing array imperfections. Firstly, we train a DOA estimation model using the source domain data. A transfer learning method is then employed to align the features between the source and target domains, thereby enhancing the deep learning model’s ability to adapt to array imperfections and improving performance in scenarios with low SNR, small snapshots, and multiple array imperfections. Moreover, in a variety of scenarios, extensive experiments have been conducted to validate the effectiveness of the proposed model. The contributions of our work can be outlined as follows:
@@
    We propose a novel DOA estimation model based on Vision Transformer, which demonstrates exceptional performance under challenging conditions, such as low SNR and small snapshot scenarios.
@@
    Due to the presence of array imperfections, the data distributions of the source and target domains differ significantly, leading to substantial performance degradation when the model is deployed in practical scenarios. To address this issue, we introduce a transfer learning algorithm to align the features between the source and target domains, enhancing the model’s performance in practical scenarios.
@@ II Problem Formulation
Figure 1: The framework of DOA-ViT and transfer learning approach proposed for array imperfections.
In this work, two models of array receiving signals are considered. The first is the ideal array signal model, where the signal received by each sensor is represented as $Be^{j\varphi}s(t)$. where $B$ denotes the sensor’s gain for the incoming signal, and $e^{j\varphi}$ accounts for the phase shift caused by the time delay of the signal propagation between the sensors. This idealized model aligns with the standard assumptions used in most DOA estimation methods.
The second signal model incorporates array imperfections to reflect real-world conditions. It considers factors such as gain and phase inconsistencies, sensor position biases, and mutual coupling. These imperfections introduce additional “array imperfection noise”, which affects the received signal and captures the complexities of practical scenarios.
@@ II-A Ideal Signal Model
where $\boldsymbol{\theta}=\left[\theta_{1},\ldots,\theta_{K}\right]$ represents the directions of arrival, $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})=\left[\boldsymbol{a}\left(\boldsymbol{p},\theta_{1}\right),\ldots,\boldsymbol{a}\left(\boldsymbol{p},\theta_{K}\right)\right]$ denotes the $M\times K$ array manifold matrix, $t=1,\ldots,T$ is the time index, and $T$ is the number of received snapshots. $\boldsymbol{s}(t)={{[{s_{1}}(t),\ldots,{s_{K}}(t)]}^{T}}$ and $\boldsymbol{n}(t)$ denote the transmitted signal and the noise vector at time index $t$, respectively. $\boldsymbol{s}(t)$ is Gaussian distributed in both its real and imaginary parts, and $\boldsymbol{n}(t)$ is additive Gaussian noise, with no correlation between any of the signals or noise. For a ULA, $\boldsymbol{a}(\boldsymbol{p},\theta_{k})$ can be expressed as:
$\boldsymbol{a}\left(\boldsymbol{p},\theta_{k}\right)=\left[1,e^{-j\frac{2\pi d}{\lambda}\sin\left(\theta_{k}\right)},\ldots,e^{-j\frac{2\pi d}{\lambda}(M-1)\sin\left(\theta_{k}\right)}\right]^{T}$ (2)
where $\lambda=c/f$ denotes the wavelength of the incident signals, $d$ is the distance between the array elements, while $c$ and $f$ represent the propagation speed and the frequency of the incident signals, respectively. Similarly, for a UCA, $\boldsymbol{a}(\boldsymbol{p},\phi_{k},\theta_{k})$ can be represented in the following form:
$\boldsymbol{a}\left(\boldsymbol{p},\phi_{k},\theta_{k}\right)=$ $\left[e^{-j\left(2\pi/\lambda\right)\cos(\phi_{1}-\theta_{k})\sin(\phi_{k})},\dots,\right.$ (3)
$\left.e^{-j\left(2\pi/\lambda\right)\cos(\phi_{M}-\theta_{k})\sin(\phi_{k})}\right]^{T}$
where $\phi_{m}=\frac{2\pi m}{M}$ denotes the azimuth of the $m$th sensor of the UCA. Based on the received signal vector discussed above, the spatial covariance matrix (SCM) is computed as follows:
[…]
where $\mathbb{E}\{\cdot\}$ denotes the expectation operator, ${{\boldsymbol{R}}_{s}}=\mathbb{E}\left\{\boldsymbol{s}(t){{\boldsymbol{s}}^{\mathrm{H}}}(t)\right\}$ represents the covariance matrix of incident signals and $\sigma_{N}^{2}$ stands for the noise power. Depending on the statistical degrees of freedom of SCM, $K\leq M$ different DOAs can be estimated precisely. However, since the number of snapshots is limited in practice, the SCM is replaced by the sampled SCM calculated via:
@@
$\widetilde{\boldsymbol{R}}_{y}$ is an unbiased estimate of ${\boldsymbol{R}}_{y}$ calculated by the received signal vector. Most existing DOA estimation algorithms focus on deriving the source DOAs $\boldsymbol{\theta}=[{{\theta}_{1}},\ldots,{{\theta}_{K}}]$ from the sampled SCM $\widetilde{\boldsymbol{R}}_{y}$, which has been proven to be the most effective approach.

### v5

# TransDOA: Calibrating Array Imperfections via Transformer-based Transfer Learning

@@ Abstract
In the proposed approach, we highlight three major contributions. First, we propose a Transformer based method for DOA estimation, which achieves excellent performance in scenarios with low signal-to-noise ratios (SNR) and limited snapshots. Second, we introduce a transfer learning framework that extends deep learning models from ideal simulation scenarios to complex real-world scenarios with array imperfections. By leveraging prior knowledge from ideal simulation data, the proposed transfer learning framework significantly improves deep learning-based DOA estimation performance in the presence of array imperfections, without the need for extensive real-world data. Finally, we incorporate visualization and evaluation metrics to assess the performance of DOA estimation algorithms, which allow for a more thorough evaluation of algorithms and further validate the proposed method. Our code can be accessed at https://github.com/zzb-nice/DOA_est_Master.
@@ I INTRODUCTION
Figure 1: Schematic illustration of environmental perception using front-mounted mmWave radar in a night-time autonomous driving scenario. The radar point clouds are projected onto the camera image for visualization.
Direction of Arrival (DOA) estimation plays a pivotal role across diverse engineering disciplines, including wireless communications, radio astronomy, radar, and sonar [1, 2, 3, 4], where precise signal localization is critical. Recently, the proliferation of accessible wireless sensing technology has expanded the scope of DOA algorithms to emerging commercial domains, such as autonomous driving and unmanned aerial systems (UAS) [5, 6, 7, 8].
Rapid advancements in engineering have imposed increasingly stringent requirements on DOA estimation algorithms. For instance, autonomous driving necessitates the tracking of fast-moving targets with small radar cross-sections [5], while non-cooperative target localization demands both high speed and precision [9]. In practice, these tasks are significantly complicated by the coexistence of challenging signal environments (e.g., low SNR and limited snapshots) and inevitable hardware impairments.
Fig. 1 illustrates the impact of signal-level and hardware-level challenges [10]. In this real-world autonomous driving scenario, their interplay leads to severe performance degradation, manifested as significant point cloud sparsity and missed detections (e.g., the sparsely represented vehicle in the red box). However, existing algorithms often fail to address these coupled constraints simultaneously. Therefore, there is a compelling need for a new algorithm that delivers robust performance against array mismatches while maintaining high accuracy in low SNR and limited snapshot regimes.
@@
Existing DOA estimation methods can be broadly categorized into model-driven and data-driven approaches. Among them, model-driven methods have been more extensively investigated due to their well-established theoretical foundations. Based on different estimation strategies, model-driven approaches can be further classified into several primary categories, including beamforming methods [11], subspace-based methods [12, 13, 14], maximum likelihood methods [15, 16], and compressed sensing (sparsity-inducing) methods [20, 17, 18, 19].
@@
A leading and well-respected family of model-driven approaches is the subspace-based methods, which achieve satisfactory performance in scenarios with high SNR and large snapshots. However, their performance relies heavily on the asymptotic accuracy of the covariance matrix, which drastically degrades under low SNR and limited snapshot conditions. Moreover, these methods struggle with color noise, and require the estimated number of signal sources to be smaller than the number of array elements, which limits their applicability.
Even if these signal-level limitations could be overcome, a more fundamental, physical challenge remains: the divergence between the idealized model and the physical hardware. Most standard model-driven methods typically rely on an ideal array manifold. However, The existence of array position biases, amplitude and phase inconsistencies, mutual coupling [21], and other array imperfections can severely distort the received data. In such cases, the mismatch between the theoretical model and the physical array often leads to significant performance degradation or even the complete failure of DOA estimation.
To mitigate these hardware impairments, various active calibration [22, 23, 24] and self-calibration [25, 26, 27, 28, 29, 30, 31, 32] techniques have been developed. Active calibration methods, while affording high accuracy through the utilization of predefined auxiliary sources, necessitate significant resource expenditure and introduce operational discontinuity. In contrast, self-calibration methods offer a more practical solution by estimating array parameters and DOAs simultaneously without auxiliary sources. However, these methods often rely heavily on initialization and struggle to handle more severe array imperfections. As noted in [25], most existing self-calibration methods are ineffective in practical scenarios with low SNR, limited snapshots, and spatially proximate sources——precisely the challenges that many DOA estimation algorithms face.
Moreover, existing calibration methods often lack versatility. Due to the complexity of modeling combined imperfections, most algorithms focus on correcting only a single type of array imperfections, such as array position biases [22, 24], amplitude-phase inconsistencies [26, 27, 28, 29], or mutual coupling [30, 31, 32]. Only a few studies have explored to address multiple simultaneous imperfections [25, 23], and these often rely on specific array geometries (e.g., ULA/UCA) or other restrictive assumptions [25]. Consequently, developing a robust framework capable of handling multiple array imperfections simultaneously, especially under low SNR and limited snapshot regimes, remains an urgent and open research problem.
In recent years, the rapid advancement of computational hardware [33] has significantly accelerated the development of deep learning technologies. Empirical studies have demonstrated the superior capability of deep learning-based methods to approach the Cramér-Rao Lower Bound (CRLB) in challenging signal regimes characterized by low SNR and limited snapshots [34, 35]. As a result, data-driven methods are considered promising tools for achieving DOA estimation in practical scenarios.
@@
Additionally, recent studies have concentrated on combining traditional DOA estimation methodologies with deep learning-based architectures to realize further advancements in performance. In 2022, the method proposed in [36] integrated subspace-based methods with deep neural networks. Specifically, a CNN was employed to reconstruct a noise-free covariance matrix, after which a classical subspace-based approach was applied for DOA estimation. Similarly, the approaches in [37, 38] combined deep learning with compressed sensing algorithms. In these works, a DNN is utilized to the extend the covariance matrix to a higher-dimensional space or to compute prior knowledge for the compressed sensing algorithm. Then a sparse reconstruction methods is applied to estimate the incident angles. Furthermore, in 2024, Shmuel et al. [39] introduced SubspaceNet, which directly treated the Root MUSIC algorithm as part of the network. Unlike previous work, SubspaceNet utilized the differentiability of polynomial equation root-finding and matrix decomposition, which allowed for the backpropagation of loss from the Root MUSIC algorithm.
Although these deep learning architectures have proven highly effective in challenging signal environments, the potential of utilizing deep learning techniques to mitigate physical array imperfections remains significantly unexplored in existing literature. While Liu et al. [35] suggested that deep learning algorithms exhibit good adaptability to array imperfections, this adaptability does not extend to all operating conditions. As will be empirically demonstrated in Fig. 12 and Fig. 14, deep learning models exhibit a pronounced sensitivity to array imperfections in challenging signal environments. Therefore, developing a robust deep learning-based approach capable of handling the simultaneous coexistence of adverse signal conditions and inevitable hardware impairments remains a critical and unresolved research challenge.
@@
As an emerging deep learning technique, transfer learning algorithm [40] is capable of leveraging knowledge acquired from a source task to enhance the performance of a distinct target task. This characteristic holds the potential to enhance algorithmic resilience in practical scenarios, particularly where challenging signal environments and inevitable hardware impairments coexist. Nonetheless, there is currently a lack of research on the application of transfer learning methods in the field of DOA estimation. To the best of our knowledge, only a few papers [41, 42, 43] have introduced transfer learning methods into DOA estimation. In [41], Wu et al. migrated a network designed for DOA estimation with multiple targets to one for fewer targets, with the primary goal of reducing the computational cost of training. Similarly, in [43], Labbaf et al. migrated a network designed for single-target DOA estimation to multiple targets. In [42], Zhang et al. applied transfer learning to enhance the network’s adaptability to array imperfections, but they only fine-tuned the network from ideal data to array imperfections-affected data, without incorporating more advanced transfer learning techniques.
@@
In this paper, we propose a Transformer-based transfer learning framework for DOA estimation, which aims to mitigate the performance degradation of data-driven approaches in practical scenarios with various array imperfections. In our approach, the data is divided into source and target domains. The source domain data consists of ideal data without array imperfections, while the target domain data represents practical scenarios containing array imperfections. Firstly, we propose a DOA estimation model, termed TransDOA, which is trained initially on the source domain data. A transfer learning method is then employed to align the features between the source and target domains, thereby significantly enhancing TransDOA’s ability to adapt to array imperfections and improving performance in scenarios with low SNR, limited snapshots, and multiple array imperfections. Moreover, in a variety of scenarios, extensive experiments have been conducted to validate the effectiveness of the proposed approach. The contributions of our work can be outlined as follows:
@@
    We propose a novel DOA estimation model, TransDOA. This model demonstrates exceptional performance under challenging conditions, such as low SNR and limited snapshot scenarios.
@@
    Due to the presence of array imperfections, the data distributions of the source and target domains differ significantly, leading to substantial performance degradation when the model is deployed in practical scenarios. To address this issue, we introduce a transfer learning algorithm to align the features between the source and target domains, enhancing TransDOA’s performance in practical scenarios.
@@
This paper is organized as follows. Section II formulates the problem. Section III presents the design of a network for DOA estimation. Section IV introduces a transfer learning scheme for the developed model. Experimental results are given in Section V. Finally, Section VI concludes the paper.
@@ II Problem Formulation
Figure 2: The framework of TransDOA and transfer learning approach proposed for array imperfections.
In this work, two models of array receiving signals are considered. The first is an ideal array signal model, while the second incorporates array imperfections to reflect real-world conditions.
@@ II-A Ideal Signal Model
where $\boldsymbol{\theta}=\left[\theta_{1},\ldots,\theta_{K}\right]$ represents the directions of arrival, $\boldsymbol{A}(\boldsymbol{p},\boldsymbol{\theta})=\left[\boldsymbol{a}\left(\boldsymbol{p},\theta_{1}\right),\ldots,\boldsymbol{a}\left(\boldsymbol{p},\theta_{K}\right)\right]$ denotes the $M\times K$ array manifold matrix, $t=1,\ldots,T$ is the time index, and $T$ is the number of received snapshots. $\boldsymbol{s}(t)={{[{s_{1}}(t),\ldots,{s_{K}}(t)]}^{T}}$ and $\boldsymbol{n}(t)$ denote the transmitted signal and the noise vector at time index $t$, respectively. The steering vectors for ULA and UCA can be expressed in the following form:
$\boldsymbol{a}(\boldsymbol{p},\theta_{k})=[1,e^{-j\frac{2\pi d}{\lambda}\sin(\theta_{k})},\ldots,e^{-j\frac{2\pi d}{\lambda}(M-1)\sin(\theta_{k})}]^{T}$ (2)
$\boldsymbol{a}\left(\boldsymbol{p},\theta_{k},\phi_{k}\right)=$ $\left[e^{-j\frac{2\pi R}{\lambda}\cos(\phi_{1}-\theta_{k})\sin(\phi_{k})},\dots,\right$. (3)
$\left.e^{-j\frac{2\pi R}{\lambda}\cos(\phi_{M}-\theta_{k})\sin(\phi_{k})}\right]^{T}$
where $\theta_{k}$ and $\phi_{k}$ denote the DOA of the $k$-th signal. $\lambda=c/f$ denotes the wavelength of the incident signals, $d$ is the distance between the array elements in ULA and $R$ is the radius of the UCA. $\phi_{m}=\frac{2\pi m}{M}$ denotes the azimuth of the $m$th sensor of the UCA. Based on the received signal discussed above, the sampled covariance matrix (SCM) is computed as follows:
@@
$\widetilde{\boldsymbol{R}}_{y}$ is an unbiased estimate of ideal covariance matrix calculated by the received signal vector. Most existing DOA estimation algorithms focus on deriving the source DOAs from the sampled SCM $\widetilde{\boldsymbol{R}}_{y}$, which has been proven to be an effective approach.